OpenStudy (anonymous):

Find the integral of x^2 +1 / x^3+3x

OpenStudy (zarkon):

this? \[\int\frac{x^2+1}{x^3+3x}dx\]

OpenStudy (anonymous):

yes

OpenStudy (zarkon):

\[u=x^3+3x\]

OpenStudy (anonymous):

and how do i do it after that ?

OpenStudy (zarkon):

\[du=(3x^2+3)dx\] \[du=3(x^2+1)dx\] \[\frac{du}{3}=(x^2+1)dx\]